OpenStudy (anonymous):

Kenny solved the equation t plus two-thirds equals one-sixth as follows: Line one shows: t plus two-thirds minus two-thirds equals one-sixth minus two-thirds. Line two shows: t equals one-sixth minus two-thirds. Line three shows: t equals one-sixth minus four-sixths. Line four shows: t equals one-half. Which statement best describes his error? He subtracted the wrong fraction on both sides of the equal sign. He subtracted when he should have added two-thirds to both sides of the equal sign. He did not rename the fractions correctly when establishing a common denominator. He did n

OpenStudy (anonymous):

@Chineseboy15

OpenStudy (anonymous):

First, let us see his work. t+2/3=1/6 (1) t+2/3-2/3=1/6-2/3 (2) t=1/6-2/3 (3) t=1/6-4/6 (4) t=1/2 He wronged in the line 4. line 4 should be t=-1/2 He subtracted 1/6 from 4/6, but to get the right answer, he must subtract 4/6 from 1/6.

OpenStudy (anonymous):

I think the right answer choice is the first one.

OpenStudy (anonymous):

@iceicebaby

OpenStudy (anonymous):

ok